Bill has twice as much money as Bob. Paul has $12 more than Bill. Together they all have $92. How much does Bob have?
Answer by rapaljer(4551)

Let x= Bob
2x=Bill
2x+12=Paul

x+2x+2x+12=92
5x+12= 92
5x=80
x=16 Bob
2x=32 Bill
2x+12= 44 = Paul

Check: 16+32+44=92 total

Dr. Robert J. Rapalje, Retired